# Enumeration of maps with tight boundaries and the Zhukovsky transformation

### v1

@@ 4 The quasi-polynomiality phenomenon
Precisely, given a positive integer $n$, let us call parity-dependent quasi-polynomial in $n$ variables a function $f$ of $n$ integer variables, such that there exists a (necessarily unique) family of polynomials $(f_{I})_{I}$ indexed by subsets $I$ of $\{1,\ldots,n\}$, such that $f(m_{1},\ldots,m_{n})=f_{I}(m_{1},\ldots,m_{n})$ whenever the $m_{i}$ with $i\in I$ are odd integers, and the $m_{i}$ with $i\notin I$ are even integers. By coefficients of $f$, we mean the collection of coefficients of the polynomials $f_{I}$, which may belong to an arbitrary ring; by total degree of $f$, we mean the maximum of the total degrees of the $f_{I}$. It is convenient to allow the value $n=0$, in which case it is understood that $f$ is just a constant element of the ring at hand.
@@
which, for $02-2g-n<$ 0, is the Euler characteristic of the moduli space $\mathcal{M}_{g,n}$ [Harer1986].
@@
Let $g,n$ be nonnegative integers such that $02-2g-n<$ 0. Then, there exists a parity-dependent quasi-polynomial $\mathfrak{T}^{(g)}_{n}$ in $n$ variables, of total degree $3g-3+n$, such that for any nonnegative integers $\ell_{1},\ldots,\ell_{n}$, we have
@@
The quasi-polynomiality phenomenon was first observed in the language of topological recursion by Norbury and Scott [NoSc13], who showed that the coefficients $\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ appearing in the series expansion (10) of $\omega^{(g)}_{n}(z_{1},\ldots,z_{n})$ are quasi-polynomials in $\ell_{1},\ldots,\ell_{n}$ which are odd in each variable. We review their result in Section 4.1: combining it with the trumpet decomposition yields (87) in the case where $n$ and all $\ell_{i}$ are non-zero.
However, the quasi-polynomiality of $\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ does not imply a priori that of $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$, as there may be pathologies when some variables are set to 00. To circumvent this problem, and show that $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ indeed coincides with an even quasi-polynomial except when all variables are set to 00, we will make use in Section 4.2 of the tight boundary insertion relations established in Section 3. The proof is by induction on the number of boundaries $n$, and we still use results from topological recursion to initialise the induction.
@@
For $n=0$, Theorem 4.1 states that for any $g\geq 2$, the generating function $F^{(g)}$ of maps of genus $g$ without boundaries reads
@@
with $\mathfrak{F}^{(g)}$ a rational function with rational coefficients of $\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}}$, $k=1,\ldots,3g-2$. As we will see below, this result is obtained by combining the property known from topological recursion that $F^{(g)}$ can be expressed in terms of the so-called moments, with an expression of the moments in terms of the derivatives of $R$ and $S$.
@@
When setting the weights $t_{1},t_{2},\ldots$ for inner faces to zero, i.e. when we consider tight maps for which every face is a tight boundary, then $\mathfrak{T}^{(g)}_{n}(\ell_{1}^{2},\ldots,\ell_{n}^{2})$ reduces to $t^{2-2g-n}$ times Norbury’s lattice count (quasi-)polynomial $N_{g,n}(\ell_{1},\ldots,\ell_{n})$ [Norbury2010]. It is shown in this reference that $N_{g,n}(0,\ldots,0)=\chi(\mathcal{M}_{g,n})$. This is consistent, in view of (87), with the fact that $T^{(g)}_{0,\ldots,0}$ vanishes when we set the weights for inner faces to zero (there are no tight maps having only boundary-vertices and no faces).
@@ 4.1 Maps without boundary-vertices
In this section, we establish the quasi-polynomiality property (87) in the case where $n$ and all $\ell_{i}$ are non-zero. We will do so by combining Theorem 2.5 with known results coming from the theory of topological recursion. One reason behind the fact that this theory is formulated in terms of $\omega^{(g)}_{n}$ instead of $W^{(g)}_{n}$ is that the former are rational functions of their $n$ variables with a simple pole structure, see for instance [Eynard2016, Section 3.3.1]. It was recognized by Norbury and Scott that this simple pole structure amounts to the following proposition:
@@
For $g\geq 0$, $n\geq 1$ and $02-2g-n<$ 0, the coefficients $\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ appearing in the expansion (10) of $\omega^{(g)}_{n}$ are of the form
@@
with $\mathfrak{T}^{(g)}_{n}$ a parity-dependent quasi-polynomial in $n$ variables of total degree $3g-3+n$.
The above proposition is a reformulation of [NoSc13, Theorem 1] [^5] Precisely, this theorem considers the expansion of $\omega^{(g)}_{n}(z_{1},\ldots,z_{n})$ around $z_{i}=0$. Here, our $\hat{\tau}_{\ell_{1},\ldots,\ell_{n}}^{(g)}$ are rather defined in terms of the expansion around $z_{i}=\infty$. But the two expansions can be related through antisymmetry properties of $\omega^{(g)}_{n}$, see (94)., in which the quantity $\mathfrak{T}^{(g)}_{n}(\ell_{1}^{2},\ldots,\ell_{n}^{2})$ is denoted by $N^{g}_{n}(\ell_{1},\ldots,\ell_{n})$. Note that we do not characterize the structure of the coefficients of $\mathfrak{T}^{(g)}_{n}$ at this stage. The relation (89) implies that (87) holds when $n$ and all $\ell_{i}$ are non-zero, by (11) and (17).
For completeness, let us sketch a proof of Proposition 4.4 based on the results available in [Eynard2016]. We start with a technical lemma.
@@
Let $P(n)$ be a polynomial of degree at most $12d+$ 1, which is equivalent to the series $F(z):=\sum_{n\geq 0}\frac{P(n)}{z^{n+1}}$ being a rational function of the form $N(z)/(z-1)^{2d+2}$, with $N$ a polynomial of degree at most $12d+$ 1. Then, the following properties are equivalent:
@@
    $N$ satisfies $N(z)=z^{2d}N(z^{-1})$, i.e. it is a self-reciprocal (palindromic) polynomial.
@@
By [StanleyEC1, Proposition 4.2.3] [^6] Short rederivation: we consider the vector space $\mathcal{L}$ of formal Laurent series of the form $\sum_{n\in\mathbb{Z}}a_{n}z^{n}$. We cannot multiply two elements of this space in general, but we can multiply an element of $\mathcal{L}$ by a polynomial in $z$, and this operation is linear. Let us consider in particular the series $\sum_{n\in\mathbb{Z}}\frac{P(n)}{z^{n+1}}\in\mathcal{L}$, and the polynomial $(z-1)^{2d+2}$. Their product vanishes since $P$ is a polynomial of degree at most $12d+$ 1 hence is annihilated by the $(2d+2)$-th power of the forward difference operator. We deduce that $(z-1)^{2d+2}\sum_{n\geq 0}\frac{P(n)}{z^{n+1}}=-(z-1)^{2d+2}\sum_{n\leq-1}\frac{P(n)}{z^{n+1}}$. The left-hand side may be interpreted as a product in the ring $\mathbb{C}((z^{-1}))$ and is thus equal to $N(z)\in\mathbb{C}[z]$. Therefore the right-hand side is also equal to $N(z)$, and we obtain an identity valid in $\mathbb{C}[[z]]$. Multiplying by the inverse of $(z-1)^{2d+2}$ in $\mathbb{C}[[z]]$ (which differs from that in $\mathbb{C}((z^{-1}))$!) and changing $n$ in $-n$, we obtain the wanted series identity., the rational function $F(z)$ admits around 00 the expansion $F(z)=\allowbreak-\sum_{n\geq 1}P(-n)z^{n-1}$ hence we have
@@
This implies the equivalence between (i) and (ii). The equivalence between (ii) and (iii) is straightforward. ∎
@@
According to [Eynard2016, Lemma 3.3.3 page 89, Theorem 3.5.1 page 121, and last equation on page 140] [^7] There seems to be a discrepancy between Theorem 3.5.1 and the equation on page 140: $\sum d_{i}$ is at most $6g-6+2n$ in the former, and $6g-4+2n$ in the latter. We assume that the former is correct since it holds for pairs of pants ($g=0,n=3$) and lids ($g=1,n=1$)., $\omega^{(g)}_{n}$ has a partial fraction decomposition of the form
@@
where the $c_{\epsilon,d}$ are some coefficients depending on the spectral curve. This implies that, in the expansion (10) of $\omega^{(g)}_{n}$, the $\hat{\tau}$ coefficients are of the form
@@
where $Q^{(g,\epsilon)}$ is a polynomial of degree at most $6g-6+3n$. It remains to check that these polynomials are indeed odd in each variable, so that we may write, for some polynomials $P^{(g,\epsilon)}$ of degree at most $3g-3+n$,
@@
For this we use the antisymmetry property of [Eynard2016, Lemma 3.3.2, page 88], which entails that
$\omega^{(g)}_{n}(z_{1},z_{2},\ldots,z_{n})=z_{1}^{-2}\omega^{(g)}_{n}(z_{1}^{-1},z_{2},\ldots,z_{n})$ (94)
and similarly for all other variables. In the expansion (91), each term of the sum over $\epsilon$ must be separately invariant under the antisymmetry mapping $f(z_{i})\mapsto z_{i}^{-2}f(z_{i}^{-1})$. Such term corresponds to the term $Q^{(g,\epsilon)}(\ell_{1},\ldots,\ell_{n})\epsilon_{1}^{\ell_{1}}\cdots\epsilon_{2}^{\ell_{2}}$ in (92) and we claim that $Q^{(g,\epsilon)}$ is odd in each of its $n$ variables. Indeed, by the transformation $z_{i}\mapsto\epsilon_{i}z_{i}$, it suffices to consider the case $\epsilon=(1,\ldots,1)$, and the claim results from Lemma 4.5 and induction over $n$. As $m_{i}=\ell_{i}^{2}$ has the same parity as $\ell_{i}$, we conclude that (89) holds with
@@
which is indeed a parity-dependent quasi-polynomial in $n$ variables of total degree $3g-3+n$. ∎
@@ 4.2 Maps with boundary-vertices
Let $g,n$ be nonnnegative integers such that $02-2g-n<$ 0. There exists an element $Q^{(g)}_{n}(x_{1},\ldots,x_{3g-2+n};u_{1},\ldots,u_{n})\in\mathbb{Q}(x_{1},\ldots,x_{3g-2+n})[u_{1},\ldots,u_{n}]$ and a constant $c(g,n)\in\mathbb{Q}$ such that, for any integers $\ell_{1},\ldots,\ell_{n}$, we have
@@
$\left.T^{(g)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}=D_{2\ell_{n+1}}\left.T^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\\ \sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}+\ell_{n+1}-1}R^{\prime}\left.T^{(g)}_{2\ell_{1},\ldots,2\ell_{i-1},2m_{i},2\ell_{i+1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}\,.$ (97)
@@
In the second line of (97), it is understood that, if $\ell_{i}=0$, then the corresponding sum over $0<m_{i}<\ell_{i}$ vanishes. In terms of the $\tau$’s, this relation rewrites as
$\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}=\frac{D_{2\ell_{n+1}}\!\!\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}}{R^{\ell_{n+1}}}\\ +\frac{R^{\prime}}{R}\sum_{i=1}^{n}\left(\ell_{i}\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\sum_{0<m_{i}<\ell_{i}}(2m_{i})\!\!\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{i-1},2m_{i},2\ell_{i+1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}\right)\,.$ (99)
@@
Here, the first term is clearly polynomial in $\ell_{1}^{2},\ldots,\ell_{n}^{2}$. Moreover, it is also a polynomial in $\ell_{n+1}^{2}$ by (79), which also entails that the coefficients are rational in $R^{(k)}/R$ for $1\leq k\leq 3g-2+(n+1)$. Finally, the second term in the last display is zero unless $\ell_{1},\ldots,\ell_{n+1}$ all vanish, in which case it is equal to $-c(g,n)\frac{\partial}{\partial t}t^{2-2g-n}=-c(g,n+1)t^{2-2g-(n+1)}$, where $c(g,n+1)=(2-2g+n)c(g,n)$. This completes the proof of the induction step: if the statement of Proposition 4.6 is true for $(g,n)$, then it is also true for $(g,n+1)$.
It remains to initialise the induction. For $g=0$, we initialise the induction at $n=3$ by using the explicit formula of [triskell, Theorem 1.1], which is the bipartite version of Theorem A.1 in Appendix A. For $g\geq 1$, we rely on consequences of the topological recursion that are described in [budd2020irreducible]. Assume first that $g\geq 2$, in which case we initialise the induction at $n=0$: in view of Remark 4.2, we only have to check that the generating function $\left.F^{(g)}\right|_{\mathrm{bip}}=\left.\tau^{(g)}_{\varnothing}\right|_{\mathrm{bip}}$ of essentially bipartite maps of genus $g$ without boundary is a rational function of $R^{(k)}/R,1\leq k\leq 3g-2$. By [budd2020irreducible, Equations (31) and (32)], there exists a polynomial $\tilde{P}_{g}$ with rational coefficients in $33g-$ 3 variables such that for $t=1$ we have
@@
where $\bar{M}_{p}$ are the so-called moments, defined at [budd2020irreducible, Equation (30)]. Using Euler’s relation, we can restore the dependence on $t$ by the formula
@@
Consequently, letting $\bar{M}_{p}^{(0)}=t\bar{M}_{p}(t_{2},tt_{4},t^{2}t_{6},\ldots)$, in accordance with [budd2020irreducible, Equation (43)], we have
@@
Finally, [budd2020irreducible, Lemma 9] shows that $\bar{M}_{0}^{(0)}=R/R^{\prime}$, while we have
@@
In the case $g=1$, we rely on [budd2020irreducible, Equation (31)], stating, in our notation, that $F^{(1)}=\ln(tR^{\prime}/R)/12$. This entails, again by the boundary insertion formula and by (79), that
@@
which is of the wanted form (96), with $c(1,1)=-1/12$. This concludes the proof of Proposition 4.6. ∎
@@
Let $g,n$ be nonnnegative integers such that $02-2g-n<$ 0. There exists a parity-dependent quasi-polynomial $\mathfrak{T}^{(g)}_{n}$ in $n$ variables, of total degree $3g-3+n$, whose coefficients are rational functions with rational coefficients of the quantities $\frac{R^{(k)}}{R}$ and $\frac{S^{(k)}}{\sqrt{R}}$ for $k=1,\ldots,3g-2+n$, and a constant $c(g,n)\in\mathbb{Q}$ such that, for any integers $\ell_{1},\ldots,\ell_{n}$, we have

### v2

@@ 4 The quasi-polynomiality phenomenon
Precisely, given a positive integer $n$, let us call parity-dependent quasi-polynomial in $n$ variables a function $f$ of $n$ integer variables, such that there exists a (necessarily unique) family of polynomials $(f_{I})_{I}$ indexed by subsets $I$ of $\{1,\ldots,n\}$, such that $f(m_{1},\ldots,m_{n})=f_{I}(m_{1},\ldots,m_{n})$ whenever the $m_{i}$ with $i\in I$ are odd integers, and the $m_{i}$ with $i\notin I$ are even integers. By coefficients of $f$, we mean the collection of coefficients of the polynomials $f_{I}$, which may belong to an arbitrary ring; by total degree of $f$, we mean the maximum of the total degrees of the $f_{I}$. It is convenient to allow the value $n=0$, in which case it is understood that $f$ is just a constant element of the ring at hand. A parity-dependent quasi-polynomial is said symmetric if it is invariant under any permutation of its variables.
@@
which, for 22gn<02-2g-n<0, is the Euler characteristic of the moduli space $\mathcal{M}_{g,n}$ [HZ86].
@@
Let $g\geq 0,n\geq 1$ be integers such that 22gn<02-2g-n<0. Then, there exists a symmetric parity-dependent quasi-polynomial $\mathfrak{T}^{(g)}_{n}$ in $n$ variables, of total degree 3g3+n3g-3+n, such that for any nonnegative integers $\ell_{1},\ldots,\ell_{n}$, we have
@@
Recall from Remark 2.10 that there is no quasi-polynomiality phenomenon for $(g,n)=(0,2)$. The quasi-polynomiality phenomenon was first observed in the language of topological recursion by Norbury and Scott [NS13], who showed that the coefficients $\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ appearing in the series expansion (10) of $\omega^{(g)}_{n}(z_{1},\ldots,z_{n})$ are quasi-polynomials in $\ell_{1},\ldots,\ell_{n}$ which are odd in each variable. We review their result in Section 4.1: combining it with the trumpet decomposition yields (88) in the case where $n$ and all $\ell_{i}$ are non-zero.
However, the quasi-polynomiality of $\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ does not imply a priori that of $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$, as there may be pathologies when some variables are set to 0. To circumvent this problem, and show that $\tau^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ indeed coincides with an even quasi-polynomial except when all variables are set to 0, we will make use in Section 4.2 of the tight boundary insertion relations established in Section 3. The proof is by induction on the number of boundaries $n$, and we still use results from topological recursion to initialise the induction.
@@
In fact, Theorem 4.1 also makes sense for $n=0$, and states that for any $g\geq 2$, the generating function $T^{(g)}_{\varnothing}=F^{(g)}$ of maps of genus $g$ without boundaries is equal to
@@
with $\mathfrak{F}^{(g)}$ a rational function with rational coefficients of $\frac{R^{(k)}}{R},\frac{S^{(k)}}{\sqrt{R}}$, $k=1,\ldots,3g-2$. As we will see below, this result is obtained by combining the property known from topological recursion that $F^{(g)}$ can be expressed in terms of the so-called moments, with an expression of the moments in terms of the derivatives of $R$ and $S$. In the case $g=1$, we no longer have rational functions but rather logarithms, see Theorem C.1.
@@
When setting the weights $t_{1},t_{2},\ldots$ for inner faces to zero, i.e. when we consider tight maps for which every face is a tight boundary, then $\mathfrak{T}^{(g)}_{n}(\ell_{1}^{2},\ldots,\ell_{n}^{2})$ reduces to $t^{2-2g-n}$ times Norbury’s lattice count (quasi-)polynomial $N_{g,n}(\ell_{1},\ldots,\ell_{n})$ [Nor10]. It is shown in this reference that $N_{g,n}(0,\ldots,0)=\chi(\mathcal{M}_{g,n})$. This is consistent, in view of (88), with the fact that $T^{(g)}_{0,\ldots,0}$ vanishes when we set the weights for inner faces to zero (there are no tight maps having only boundary-vertices and no faces).
@@ 4.1 Maps without boundary-vertices
In this section, we establish the quasi-polynomiality property (88) in the case where $n$ and all $\ell_{i}$ are non-zero. We will do so by combining Theorem 2.5 with known results coming from the theory of topological recursion. One reason behind the fact that this theory is formulated in terms of $\omega^{(g)}_{n}$ instead of $W^{(g)}_{n}$ is that the former are rational functions of their $n$ variables with a simple pole structure, see for instance [Eyn16, Section 3.3.1]. It was recognized by Norbury and Scott that this simple pole structure amounts to the following proposition:
@@
For $g\geq 0$, $n\geq 1$ and 22gn<02-2g-n<0, the coefficients $\hat{\tau}^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ appearing in the expansion (10) of $\omega^{(g)}_{n}$ are of the form
@@
with $\mathfrak{T}^{(g)}_{n}$ a parity-dependent quasi-polynomial in $n$ variables of total degree 3g3+n3g-3+n.
The above proposition is a reformulation of [NS13, Theorem 1] [^6] Precisely, this theorem considers the expansion of $\omega^{(g)}_{n}(z_{1},\ldots,z_{n})$ around $z_{i}=0$. Here, our $\hat{\tau}_{\ell_{1},\ldots,\ell_{n}}^{(g)}$ are rather defined in terms of the expansion around $z_{i}=\infty$. But the two expansions can be related through antisymmetry properties of $\omega^{(g)}_{n}$, see (96)., in which the quantity $\mathfrak{T}^{(g)}_{n}(\ell_{1}^{2},\ldots,\ell_{n}^{2})$ is denoted by $N^{g}_{n}(\ell_{1},\ldots,\ell_{n})$. Note that we do not characterize the structure of the coefficients of $\mathfrak{T}^{(g)}_{n}$ at this stage. The relation (90) implies that (88) holds when $n$ and all $\ell_{i}$ are non-zero, by (11) and (17).
For completeness, let us sketch a proof of Proposition 4.4 based on the results available in [Eyn16]. We start with a technical lemma.
@@
Let $P(n)$ be a polynomial of degree at most $2d+1$, which is equivalent to the series $F(z):=\sum_{n\geq 0}\frac{P(n)}{z^{n+1}}$ being a rational function of the form $N(z)/(z-1)^{2d+2}$, with $N$ a polynomial of degree at most $2d+1$. Then, the following properties are equivalent:
@@
    NN satisfies $N(z)=z^{2d}N(z^{-1})$, i.e. it is a self-reciprocal (palindromic) polynomial.
@@
By [Sta12, Proposition 4.2.3] [^7] Short rederivation: we consider the vector space $\mathcal{L}$ of formal Laurent series of the form $\sum_{n\in\mathbb{Z}}a_{n}z^{n}$. We cannot multiply two elements of this space in general, but we can multiply an element of $\mathcal{L}$ by a polynomial in $z$, and this operation is linear. Let us consider in particular the series $\sum_{n\in\mathbb{Z}}\frac{P(n)}{z^{n+1}}\in\mathcal{L}$, and the polynomial $(z-1)^{2d+2}$. Their product vanishes since $P$ is a polynomial of degree at most $2d+1$; hence, it is annihilated by the $(2d+2)$-th power of the forward difference operator. We deduce that $(z-1)^{2d+2}\sum_{n\geq 0}\frac{P(n)}{z^{n+1}}=-(z-1)^{2d+2}\sum_{n\leq-1}\frac{P(n)}{z^{n+1}}$. The left-hand side may be interpreted as a product in the ring $\mathbb{C}((z^{-1}))$ and is thus equal to $N(z)\in\mathbb{C}[z]$. Therefore the right-hand side is also equal to $N(z)$, and we obtain an identity valid in $\mathbb{C}[[z]]$. Multiplying by the inverse of $(z-1)^{2d+2}$ in $\mathbb{C}[[z]]$ (which differs from that in $\mathbb{C}((z^{-1}))$!) and changing $n$ in n-n, we obtain the wanted series identity., the rational function $F(z)$ admits around 0 the expansion $F(z)=\allowbreak-\sum_{n\geq 1}P(-n)z^{n-1}$ hence we have
@@
This implies the equivalence between (i) and (ii): indeed, if $P$ is odd, then $z^{-2}F(z^{-1})=\sum_{n\geq 1}z^{-n-1}P(n)=F(z)$, since $P(0)=0$, and conversely, if $F(z)=z^{-2}F(z^{-1})$ then by comparing the expansions, we obtain $P(0)=0$ and P(n)=P(n)-P(-n)=P(n) for every $n\geq 1$, implying that $P$ is odd. The equivalence between (ii) and (iii) is straightforward. ∎
@@
According to [Eyn16, Lemma 3.3.3 page 89, Theorem 3.5.1 page 121, and last equation on page 140] [^8] There seems to be a discrepancy between Theorem 3.5.1 and the equation on page 140: $\sum d_{i}$ is at most 6g6+2n6g-6+2n in the former, and 6g4+2n6g-4+2n in the latter. We assume that the former is correct since it holds for pairs of pants ($g=0,n=3$) and lids ($g=1,n=1$)., $\omega^{(g)}_{n}$ has a partial fraction decomposition of the form
@@
where the $c_{\epsilon,d}$ are some coefficients depending on the spectral curve. By the negative binomial formula, we have
$\frac{1}{(z-\epsilon)^{d+2}}=\sum_{\ell\geq 1}\binom{\ell}{d+1}\frac{\epsilon^{\ell-d-1}}{z^{\ell+1}}.$ (93)
Viewing $\binom{\ell}{d+1}$ as a polynomial in $\ell$, this implies that, in the expansion (10) of $\omega^{(g)}_{n}$, the $\hat{\tau}$ coefficients are of the form
@@
where $Q^{(g,\epsilon)}$ is a polynomial of degree at most 6g6+3n6g-6+3n. It remains to check that these polynomials are indeed odd in each variable, so that we may write, for some polynomials $P^{(g,\epsilon)}$ of degree at most 3g3+n3g-3+n,
@@
Thanks to the symmetry in $\ell_{1},\ldots,\ell_{n}$, it suffices to check that $Q^{(g,\epsilon)}$ is odd in $\ell_{1}$. For this we use the antisymmetry property of [Eyn16, Lemma 3.3.2, page 88], which entails that
$\omega^{(g)}_{n}(z_{1},z_{2},\ldots,z_{n})=z_{1}^{-2}\omega^{(g)}_{n}(z_{1}^{-1},z_{2},\ldots,z_{n}).$ (96)
In the expansion (92), each term of the sum over $\epsilon$ must be separately invariant under the antisymmetry mapping $f(z_{1})\mapsto z_{1}^{-2}f(z_{1}^{-1})$, since it corresponds to the contribution to $\omega^{(g)}_{n}$ having poles at $z_{1}=\epsilon_{1},\ldots,z_{n}=\epsilon_{n}$. Such term corresponds to the term $Q^{(g,\epsilon)}(\ell_{1},\ldots,\ell_{n})\epsilon_{1}^{\ell_{1}}\cdots\epsilon_{2}^{\ell_{2}}$ in (94) and we claim that $Q^{(g,\epsilon)}$ is odd in $\ell_{1}$. By the transformation $z_{i}\mapsto\epsilon_{i}z_{i}$, it suffices to consider the contribution of $\epsilon=(1,\ldots,1)$ to $\omega^{(g)}_{n}$, namely
$F(z_{1}):=\sum_{\begin{subarray}{c}d\in\mathbb{N}^{n}\\ d_{1}+\cdots+d_{n}\leq 6g-6+2n\end{subarray}}\frac{c_{(1,\ldots,1),d}}{(z_{1}-1)^{d_{1}+2}\cdots(z_{n}-1)^{d_{n}+2}},$ (97)
satisfying $F(z_{1})=z_{1}^{-2}F(z_{1}^{-1})$. By Lemma 4.5, it expands as $\sum\limits_{\ell_{1}\geq 1}\frac{P(\ell_{1})}{z_{1}^{\ell_{1}+1}}$ with $P$ an odd polynomial in $\ell_{1}$, which implicitly depends on the variables $z_{2},\ldots,z_{n}$. Further expanding in these extra variables, the coefficient of $\frac{1}{z_{2}^{\ell_{2}+1}\cdots z_{n}^{\ell_{n}+1}}$, which is nothing but $Q^{(g,(1,\ldots,1))}(\ell_{1},\ldots,\ell_{n})$, must be odd in $\ell_{1}$ as wanted, hence the oddness property is established. As $m_{i}=\ell_{i}^{2}$ has the same parity as $\ell_{i}$, we conclude that (90) holds with
@@
which is indeed a parity-dependent quasi-polynomial in $n$ variables of total degree 3g3+n3g-3+n. ∎
@@ 4.2 Maps with boundary-vertices
Let $g,n$ be nonnnegative integers such that 22gn<02-2g-n<0. There exist $Q^{(g)}_{n}(x_{1},\ldots,x_{3g-2+n};u_{1},\ldots,u_{n})\in\mathbb{Q}(x_{1},\ldots,x_{3g-2+n})[u_{1},\ldots,u_{n}]$ and a constant $c(g,n)\in\mathbb{Q}$ such that, for any integers $\ell_{1},\ldots,\ell_{n}$, we have
@@
$\left.T^{(g)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}=D_{2\ell_{n+1}}\left.T^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\\ \sum_{i=1}^{n}\sum_{0<m_{i}<\ell_{i}}(2m_{i})R^{\ell_{i}-m_{i}+\ell_{n+1}-1}R^{\prime}\left.T^{(g)}_{2\ell_{1},\ldots,2m_{i},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}\,.$ (100)
@@
In the second line of (100), it is understood that, if $\ell_{i}=0$, then the corresponding sum over $0<m_{i}<\ell_{i}$ vanishes. In terms of the $\tau$’s, this relation is rewritten as
$\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n},2\ell_{n+1}}\right|_{\mathrm{bip}}=\frac{D_{2\ell_{n+1}}\!\!\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}}{R^{\ell_{n+1}}}\\ +\frac{R^{\prime}}{R}\sum_{i=1}^{n}\left(\ell_{i}\left.\tau^{(g)}_{2\ell_{1},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}+\sum_{0<m_{i}<\ell_{i}}(2m_{i})\!\!\left.\tau^{(g)}_{2\ell_{1},\ldots,2m_{i},\ldots,2\ell_{n}}\right|_{\mathrm{bip}}\right)\,.$ (102)
@@
Here, the first term is clearly polynomial in $\ell_{1}^{2},\ldots,\ell_{n}^{2}$. Moreover, it is also a polynomial in $\ell_{n+1}^{2}$ by (80), which also entails that the coefficients are rational in $R^{(k)}/R$ for $1\leq k\leq 3g-2+(n+1)$. Finally, the second term in the last display is zero unless $\ell_{1},\ldots,\ell_{n+1}$ all vanish, in which case it is equal to $c(g,n)\frac{\partial}{\partial t}t^{2-2g-n}=-c(g,n+1)t^{2-2g-(n+1)}$, where c(g,n+1)=(22g+n)c(g,n)c(g,n+1)=(2-2g+n)c(g,n). This completes the proof of the induction step: if the statement of Proposition 4.6 is true for $(g,n)$, then it is also true for $(g,n+1)$.
It remains to initialise the induction. For $g=0$, we initialise the induction at $n=3$ by using the explicit formula of [BGM22, Theorem 1.1], which is the bipartite version of Theorem A.1 in Appendix A. For $g\geq 1$, we rely on the consequences of the topological recursion that are described in [Bud22]. Assume first that $g\geq 2$, in which case we initialise the induction at $n=0$: in view of Remark 4.2, we only have to check that the generating function $\left.F^{(g)}\right|_{\mathrm{bip}}=\left.\tau^{(g)}_{\varnothing}\right|_{\mathrm{bip}}$ of essentially bipartite maps of genus $g$ without boundary is a rational function of $R^{(k)}/R,1\leq k\leq 3g-2$. By [Bud22, Equations (31) and (32)], there exists a polynomial $\tilde{P}_{g}$ with rational coefficients in 3g33g-3 variables such that for $t=1$ we have
@@
where $\bar{M}_{p}$ are the so-called moments, defined at [Bud22, Equation (30)]. Using Euler’s relation, we can restore the dependence on $t$ by the formula
@@
Consequently, letting $\bar{M}_{p}^{(0)}=t\bar{M}_{p}(t_{2},tt_{4},t^{2}t_{6},\ldots)$, in accordance with [Bud22, Equation (43)], we have
@@
Finally, [Bud22, Lemma 9] shows that $\bar{M}_{0}^{(0)}=R/R^{\prime}$, while we have
@@
In the case $g=1$, we rely on [Bud22, Equation (31)], stating, in our notation, that $F^{(1)}=\ln(tR^{\prime}/R)/12$. This entails, again by the boundary insertion formula and by (80), that
@@
which is of the wanted form (99), with c(1,1)=1/12c(1,1)=-1/12. This concludes the proof of Proposition 4.6. ∎
@@
Let $g,n$ be nonnnegative integers such that 22gn<02-2g-n<0. There exist a parity-dependent quasi-polynomial $\mathfrak{T}^{(g)}_{n}$ in $n$ variables, whose coefficients are rational functions with rational coefficients of the quantities $\frac{R^{(k)}}{R}$ and $\frac{S^{(k)}}{\sqrt{R}}$ for $k=1,\ldots,3g-2+n$, and a constant $c(g,n)\in\mathbb{Q}$ such that, for any integers $\ell_{1},\ldots,\ell_{n}$, we have